A solution of antifreeze contains 145 g of ethylene glycol (C2H6O2) per 550. gram of water....

  1. A solution of antifreeze contains 145 g of ethylene glycol (C2H6O2) per 550. gram of water. What is the freezing point of this solution? (Kf = 1.86 oC/m)
Answer #1

The freezing point of solution can be calculated as;

ΔTF = KF · b · i

again ,ΔTF = i X KF X (W2 X 1000) / (M2 X W1) ...................(1)

Where

i = van 't Hoff factor = 1

b =molality = (W2/M2) / (W1/1000)

W2 = weight of ethylene glycol =145 g

W1 = Weight of water = 550 g

M2 = Moler mass of ethylene glycol = 62.07 g/mol

KF = cryoscopic constant = 1.86 deg C/m =  = 1.86 K·kg/mol

Putting the value in equation (1) we get

ΔTF = (1) X (1.86 X 145 X 1000 ) / (62.07 X 550) = 269700 / 34138.5 = 7.9 K

Again,

ΔTF = TF (pure solvent) − TF (solution) ...............................(2)

Here  TF (pure solvent) 273.15

TF (solution)= TF (pure solvent) − ΔTF = 273.15 - 7.9 = 265.25 K = - 7.9 deg C

Freezing point of solution is - 7.9 deg C.